[Hallc_running] [New Logentry] Status update

mack at jlab.org mack at jlab.org
Fri Feb 23 10:10:02 EST 2018

Logentry Text:
Sign up for shifts! 

While the MPS bcm in Hall C was tuned, an opportunistic RF recovery took place. When beam came back, there was a higher current attempt, then a return to lower currents to finish a solid target kinematic setting. Then we got knocked offline due to a radiation monitor trip in the truck ramp. The offending equipment has been replaced.  Data taking resumed with currents as high as 60muA.  A so-called "high current test" is harder and more ambiguous than it sounds. With a baseline of ~8 trips/hour, an hour of beam at a given current can only tell you at high confidence if the trip rate got much better ( 2.3 or below), or got much worse (increased to 13.7 or above). We can chase our tails basing decisions on 1sigma fluctuations. Earlier in the evening, I saw no believable correlation between current and trip rate. But later it appeared that the trip rate was unacceptably high at higher currents, so owl shift ran at 40-45 muA depending on the target. The PD's plan is to adiabatically push 
 for higher currents today while experts are available, then coast thru the weekend. 
 There was several items of major to minor concern in the last day:  Q2 ramped itself down with no interlock flags, but magnet whisperer JP Chen got it running again (https://logbooks.jlab.org/entry/3536638 ); replay indicated an HMS DC ringing alarm, suggesting that this issue occasionally pops up even at the present 5.5V threshold ( https://logbooks.jlab.org/entry/3536358 );  and a pedestal shifted out of spec in an HMS calorimeter channel ( https://logbooks.jlab.org/entry/3536965 ). 

Special notices in effect for shift crews:
Don't go higher than 610A with HMS Q2 https://logbooks.jlab.org/entry/3535714
Instructions for LH2 <--> LD2 switches by TO : https://logbooks.jlab.org/entry/3534487
Maximum current for Gaskell's bor-on-ing targets is 40 muA: https://logbooks.jlab.org/entry/3535297
Any missing reference time messages should not be ignored: https://logbooks.jlab.org/entry/3535289

Good event definition tweak for hodo effi determination: https://logbooks.jlab.org/entry/3536418  and  https://logbooks.jlab.org/entry/3536031
Fix of cryomodule vacuum improves coolant overhead (or something)  https://logbooks.jlab.org/entry/3536612       
MPS bcm was tuned better but the jury's still out on the result: https://logbooks.jlab.org/entry/3536478  , and https://logbooks.jlab.org/entry/3536445 , and https://logbooks.jlab.org/entry/3536550

Recent calibration/analysis updates: 
Deepak's hms dc calibration: https://logbooks.jlab.org/entry/3536825
Debitwa got recalibrated:  https://hallcweb.jlab.org/doc-private/ShowDocument?docid=958 

We will meet in 200C at 1:30 pm (second floor of the counting house).

Link: https://bluejeans.com/543903672
Connecting directly from a room system?
1) Dial: or bjn.vc
2) Enter Meeting ID: 543 903 672
Just want to dial in on your phone?
1) Direct-dial with my iPhone or
+1.408.740.7256 (United States)
+1.888.240.2560 (US Toll Free)
+1.408.317.9253 (Alternate number)
(Global Numbers)
2) Enter Meeting ID: 543 903 672
3) Press #


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3537081
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180223/38fca5b9/attachment.html>

More information about the Hallc_running mailing list